Macro Econ

ECF110-FPD-1-2020-2 (1).docx


The Employment Act, Chapter 268 of the laws of Zambia, sets out the minimum contractual age as 16 years old and the normal retirement age is 60 years old. You have been provided with the following information about the population of Zambia and its labour force.

 

Age

Have a fulltime Job

Have a part-time job

Don’t have a job but are looking for one

Don’t have a Job and are not looking for one

Total

<10

150

150

11-15

135

135

16-20

120

30

75

80

305

21-30

20

50

25

20

115

31-50

80

100

55

40

275

51-60

250

50

150

80

530

>60

50

40

30

170

290

Total

520

270

335

675

1800

 

(i)             What is the size of the labor force in this economy?                                  [5 Marks]

(ii)           Calculate the Labor Force Participation Rate for this economy. Report as a percentage to four decimal places.                                                          [2 Marks]

(iii)         Calculate the Unemployment Rate for this economy. Report as a percentage to two decimal places.                                                                                  [3 Marks]

(iv)         What age group has the highest level of unemployment? Report as a percentage to two decimal places.                                                                              [3 Marks]
Answer #1

(i) The total labor force of an economy is the total number of people in the working age which is all the people above 15 years of age and below the retirement age of 60.

So the total labor force for this economy will be the sum total of people who are above the age of 15 years and below 60 years which means they are capable of working.

Labor force = Total population -( people below the age of 16 years + people above 60 years)

Labor force = 1800 - (150 + 135 + 290)

Labor force = 1800 - 575

Labor force = 1225

So the labor force of this economy is 1225 people.

(ii) The labor force participation is the ratio of people who population which is working (whether full time or half time) plus the population which is looking for work to the total labor force. Please note that people who are part of labor force and don't have work but not looking for work are not included in labor force participation.

Here total number of people who are working full time and are below the age of 60 is 520 - 50 = 470 and people working half time and belowtthe age of 60 years is 270 - 40 = 230 and people looking for work and are below 60 years is 335 - 30 = 305

So the total is is equal to 470 + 230 + 305 = 1005

So the labor force participation = 1005/1225

Labor force participation = 0.820408

Now to convert it into percentage we need to multiply it with 100.

Labor force participation = 0.820408×100

Labor force participation = 82.0408%

So the labor force participation is 82.0408%

(iii) The unemployment rate is simply the total number of people who don't have work and who are looking for work, and they are also below 60 years which we calculated above as 305 divided by the labor force participants which are either working or looking for work which we calculated as 1005.

So the unemployment rate = 305/1005

Unemployment rate = 0.303482

Again to convert it into percentage we just need to multiply it by 100.

Unemployment rate = 0.303482×100

Unemployment rate = 30.3482%

So the unemployment rate for this economy is 30.3482%.

​​​

(iv) Let's calculate the unemployment rate for each age group.

Age group of 16-20.

Unemployment rate = 75/(120+30+75) × 100

Unemployment rate = 75/225 × 100

Unemployment rate = 33.3333%

For age group 21-30

Unemployment rate = 25/(20+50+25) × 100

Unemployment rate = 25/95 × 100

Unemployment rate = 26.3157%

For age group 31-50

Unemployment rate = 55/(80+100+55) × 100

Unemployment rate = 55/235 × 100

Unemployment rate = 23.4042%

For the age group 51-60.

The unemployment rate

= 150/(250+50+150) × 100

= 150/450 × 100

= 33.3333%

So the age group which has the highest unemployment rate is age group 16-20 and age group 51-60 which both has unemployment rate of 33.33%.

​​​​​​